Give an example of a function of two variables whose limit at (0, 0) does not exist, that is


lim(x,y)→(0,0) f(x, y) does not exist. Explain also why the limit does not exist.

1
Expert's answer
2021-11-24T10:25:04-0500

"lim_{(x,y)\to(0,0)}\frac{x²-y²}{x²+y²}"

This is an example of a function of two variables whose limit at (0, 0) does not exist.

The reason of non existence of the limit is discussed below.

Let us consider y = mx. So y"\to0" as x "\to 0" for any arbitrary m.

So the limit becomes

"lim_{x\to0}\frac{x²-m²x²}{x²+m²x²}"

= "lim_{x\to0}\frac{1-m²}{1+m²}"

= "\frac{1-m²}{1+m²}"

So the value of the limit varies as the value of m varies.

Hence the limit does not exist